OpenStudy (anonymous):

can someone help me with this problem: find an equation for the situation that you have 40 animals(cows and chickens) and 100 legs

OpenStudy (anonymous):

do you know how many chickens and how many cows?

OpenStudy (anonymous):

simple yes or no will do...

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

sorry! the total number or cows +chickens is 40 but i dont know how many of each

OpenStudy (anonymous):

ok.... since we don't know, let's let: x be the number of chickens, and y be the number of cows... BUT you do know this: x + y = 40, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now.. assuming these farm animals didn't come from Cherynoble, each chicken has 2 legs, and each cow has 4 legs.... correct?

OpenStudy (anonymous):

yes so wouldnt that be 2x+4y=100?

OpenStudy (anonymous):

good.... so you have the system of equations : x + y = 40 2x + 4y = 100 do you know how to solve the system?

OpenStudy (anonymous):

no thats the problem i dont know how to put the two equations together

OpenStudy (anonymous):

there are a few ways to solve systems of equations.... of these two, which one do you know? 1 - substitution 2 - elimination

OpenStudy (anonymous):

oh you mean like i make both equal to x and then put them together?

OpenStudy (anonymous):

yes you can do that too.... :)

OpenStudy (anonymous):

well what were you thinking about doing? can you help me with this part too. i dont want to mess it up

OpenStudy (anonymous):

substituting.... the first equation into the second equation....

OpenStudy (anonymous):

actually elimination would be faster...

OpenStudy (anonymous):

ok how would i do that? i really dont remember... im sorry

OpenStudy (anonymous):

i'll do substitution then....

OpenStudy (anonymous):

i dont remember that either but whichever you like better

OpenStudy (anonymous):

in the first equation, solve for y... x + y = 40 ---> y = 40 - x now let's replace the y in the second equation with this expression for y... 2x + 4y = 100 2x + 4(40 - x) = 100 by doing the substitution, we now have 1 equation with 1 variable.... can you solve for x now?

OpenStudy (anonymous):

ya i think i can... thank you so much for all your help

OpenStudy (anonymous):

yw...:)